OpenStudy (anonymous):

Which improper fraction is equivalent to 2 3/9

OpenStudy (anonymous):

21/9

OpenStudy (anonymous):

21/9

OpenStudy (anonymous):

that's correct

OpenStudy (anonymous):

Multiply the denominator (9) times the whole number (2) and add the numerator (3). 9 * 2 + 3 = 18 + 3 = 21 All over the denominator (9) So that's how you get 21/9